Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

When 2 people are selected at random from a group of 4 femal

Expert replies
by Max@Math Revolution » Thu Sep 27, 2018 11:34 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

[Math Revolution GMAT math practice question]

When 2 people are selected at random from a group of 4 females and 4 males, what is the probability that at least one female is selected?

A. 5/14
B. 7/14
C. 9/14
D. 11/14
E. 13/14
Join the discussion
Source: — Problem Solving |

by swerve » Fri Sep 28, 2018 8:39 am
Total possibilities 8c2 = (8*7)/2 = 28.

Now, since we are interested in probability at least one is female, we can find the probability that all 2 are males 4c2 = (4*3)/2 = 6.

Now 6/28 = 3/14.

The probability of at least one female = 1 - (3/14) = 11/14.
Join the discussion

by Max@Math Revolution » Sun Sep 30, 2018 9:46 pm
=>
The probability that at least one female is selected is equal to 1 minus the probability that two males are selected.
There are 4C2 ways of selecting 2 males from the four males and 8C2 ways of selecting 2 people from the group of 8. Therefore, the probability of selecting two males from the group is 4C2/8C2, and the probability of selecting at least one female is
1 – 4C2/8C2 = 1 – { (4*3)/(1*2) } / { (8*7)/(1*2) } = 1 – (4*3)/(8*7) = 1 – 3/14 = 11/14

Therefore, the answer is D.
Answer: D
Join the discussion

by Scott@TargetTestPrep » Sat Oct 13, 2018 5:33 pm
Max@Math Revolution wrote:
When 2 people are selected at random from a group of 4 females and 4 males, what is the probability that at least one female is selected?

A. 5/14
B. 7/14
C. 9/14
D. 11/14
E. 13/14
We can use the formula:

P(at least one female is selected) = 1 - P(no females are selected)

P(no females are selected) = 4/8 x 3/7 = 1/2 x 3/7 = 3/14

P(at least one female is selected) = 1 - 3/14 = 11/14.

Alternate Solution:

Without any restriction, there are 8C2 = 8!/(2!*6!) = 28 different ways to choose two people from a group of eight.

We use the following formula:

#ways to choose two people = #ways where at least one female is selected + #ways where no female is selected

Notice that since there are 4 men, there are 4C2 = 4!/(2!*2!) = 6 ways to choose two people where no female is selected. Thus, 28 - 6 = 22 of the choices include at least one female. Thus, the probability that at least one female is selected is 22/28 = 11/14.

Answer: D

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion